OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−6, 4) and (2, 0). y − 4 = −one half(x + 6) y − 4 = 2(x + 6) y + 6 = −one half(x − 4) y + 6 = 2(x − 4)

OpenStudy (anonymous):

@satellite73 Will you help me?

OpenStudy (nubeer):

first you have to find the gradient.. do you how to find the gradient when two coordinates are given?

OpenStudy (aravindg):

@ahoward79 it seems you are trying to cheat in a test/quiz ?

OpenStudy (anonymous):

First Find the Slope

OpenStudy (anonymous):

then use Point Slope Form y-y1 = m (x-x1)

OpenStudy (aravindg):

our previous question was exactly similiar ,only difference the points changed

OpenStudy (anonymous):

When I tried to find the slope I got -4/8

OpenStudy (anonymous):

No actually the last one already had a slope.

OpenStudy (aravindg):

did you tell me that our answer was shown wrong or something ?

OpenStudy (anonymous):

Ok can someone actually help me?

OpenStudy (nubeer):

ok you got -4/8 which is same as -1/2 right?

OpenStudy (aravindg):

well @ahoward79 if you are actually here to learn we will be happy to help , but the thing is these questions are very similiar to each other .There is no chance that you get stuck after doing a detailed discussion on a similiar question .

OpenStudy (aravindg):

I answered the same type of questions for you more than 4 times..i am still surprised why you having doubts like these

OpenStudy (anonymous):

Ok yes @nubeer So what do I do next?

OpenStudy (anonymous):

Ok well since you arent helping you can leave my question please, thank you.

OpenStudy (aravindg):

TIP: First learn the basic formula of equation of lines thoroughly ,then attempt practise questions,be careful about signs ..thats all what is required!!

OpenStudy (aravindg):

it will do a lot more help than you asking desperately here to get the answer

OpenStudy (nubeer):

ok as Yahoo mentioned the formula above y-y1=m(x-x1) now you have m =-1/2 and y1 and x1.. so now just plug the values in formula u will get the answer.

OpenStudy (anonymous):

Which values are y1 and x1? -6 and 2?

OpenStudy (anonymous):

Nvm I figured it out its A. Thank you soo much! :)

OpenStudy (aravindg):

are you ready to learn ? if you promise me that you will study these formulas thoroughly and try your level best before posting them here ..i can help you :)

OpenStudy (anonymous):

I dont need anymore help Im ready to take my exam now.

OpenStudy (aravindg):

all the best then !

OpenStudy (nubeer):

lol best of luck.. and sorry seems to be i am just getting notifications a bit late so missed it..